# QGFace: Quality-Guided Joint Training For Mixed-Quality Face Recognition

### v1

@@ III-B2 Contrastive learning for LQ features
Although classification-based methods achieve SOTA performances on HQ face recognition, they are not suitable for learning LQ images. Fig. 2 visualizes the GST of AdaFace. It is obvious that LQ images gain more attentions during training. However, due to the low qualities, they lack the discriminative information for matching with thousands of abstractive identity features. To alleviate this problem, we apply contrastive learning on LQ features, which is based on the similarity between image instances and is more practical for learning the LQ data.
@@
To comprehensively learn the LQ images and compose the positive contrastive pairs, we apply data augmentation on the HQ training images. Different from the popular contrastive methods which apply data augmentation on the images twice to compose the positive pairs, we only augment each image once to compose the pair with the original image. Also, we stop the gradients of HQ images during training to keep the qualities of their features. Let $(\boldsymbol{x}^{q},\boldsymbol{x}^{k_{+}})$ denote the augmented image and the original image respectively, and $(\boldsymbol{h}^{q},\boldsymbol{h}^{k_{+}})$ denote their features which compose a positive pair in contrastive learning.
Different augmentation methods are used including scaling (down-sampling to a lower resolution and then up-sampling to the original size), crop & random resize, rotation, and color distortion. AdaFace introduces hard data augmentations with low possibilities including cropping 20% of the original image, which make the augmented images unrecognizable. This might be acceptable when little emphasis is put on these data, but is not suitable in our situation where all LQ features are used for learning. We increase the possibility of data augmentation with lower-level distortion. In addition, we follow [6] to add JPEG compression artifacts to make the augmented images more realistic.
Finally, to make the above two learning paradigms better cooperate with each other, we employ the Supervised Contrastive Mask (SCM) introduced in CoReFace [32]. The pair-guided contrastive learning may push two features away as they are not generated from the same image, while the classification method tries to pull them together as they are from the same identity. To address this conflict, SCM removes all samples with the same identity of the given image from the negative feature pool. Let $\theta_{qk}$, $y_{+}$ denote the angle between feature pairs and the label of the given image respectively, and $Q$, $k_{j}$ denote the number of negative samples and the $j^{th}$ feature among them. With a scale parameter $s$, the contrastive loss can be formulated as
@@ III-C Proxy-Updated Real-Time Queue
Self-supervised contrastive learning usually needs a large comparison pool to serve as the negative candidates and push the training samples away from them. MoCo [17] defines the self-supervised contrastive learning as the comparison with a queue. A slowly progressing momentum encoder is proposed to produce keys and save them in a queue. The queries are produced by the training encoder. To make the positive and the negative keys comparable, all of them are produced by the momentum encoder. However, this operation prevents the training encoder from learning the LQ and the HQ image simultaneously as one of them is sent into a gradient-free encoder. Furthermore, as the momentum encoder is designed to be stable during large training intervals, this could lose track of the status of the encoder in joint training and produce whitened features.
To relieve the above problems, we propose a novel proxy-updated queue for contrastive learning, which is illustrated in Fig. 3. We take the genuine features encoded by the training model as the positive pairs. Since the training target of the cross-entropy softmax loss could constrain the corresponding feature-proxy pair to be similar, we update the negative features with their related proxies. Thus, we build a feature queue and a proxy queue. Specifically, we first encode all the original images and the augmented ones. Then, we update the feature queue and the proxy queue with the above features and their corresponding identity proxies. As the features in queue could be generated from several iterations ago, we update them to be similar to the features generated by the training encoder. They are added with the difference between the proxies in queue and the proxies in training. To provide a mixed-quality queue, we take both $\boldsymbol{h}^{q}$ and $\boldsymbol{h}^{k}$ to update it. With SCM, the comparisons between the identical features will not be introduced.
The cross-entropy softmax loss calculates the similarity between a feature and all of the identity proxies. To provide a comparative magnitude of the pair relationship supervision, we set the size of the queue $Q$ as the number of the identities $n$. Thus, the number of pair comparison used for contrastive learning keeps large in training. This enables the contrastive learning to discriminate among more images of mixed qualities for more effective learning.
@@ IV Experiments
(a) VGGFace2 dataset
(b) High-quality dataset
(c) Mixed-quality dataset
(d) Low-quality dataset
Figure 4: (a) Examples from the training dataset VGGFace2. Each row represents an identity. Identities are randomly selected from the dataset. (b) Examples from the high-quality dataset CALFW. The positive and the negative pairs are bounded with green and red. (c) Examples from the mixed-quality dataset IJB-B. (d) Examples from the low-quality dataset Tinyface. These face crops are generated from real images and very challenging.
TABLE I: Ablations of our proposed framework. “HQ Avg.” denotes the average performance on the five HQ datasets LFW, AgeDB, CFP-FP, CALFW, and CPLFW. Rank-1 accuracy is reported for SCface and Tinyface. TAR@FAR=$10^{-4}$ is reported for IJB-B dataset.
Setting Augmentation Quality Contrastive SCface Tinyface IJB-B HQ Avg.
@@ IV-A Datasets and Implementation Details
Datasets. We take VGGFace2 [33] as the training dataset. We evaluate our approach on three categories of datasets for different tasks. The first category is the HQ datasets including the widely-used LFW [34], AgeDB [35], CFP-FP [36], CPLFW [37], and CALFW [38]. The second category is the LQ datasets including SCface dataset [12] and Tinyface dataset [23] which are usually used for the cross-resolution face recognition and the LQ face recognition respectively. SCface contains 130 identities captured by five surveillance cameras from 3 distances: 4.2m (d1), 2.6m (d2), and 1.0m (d3) for each identity. At each distance, the LQ images need to be matched with the HQ gallery images of the corresponding identities. Tinyface contains 5K identities with 153.5K distracting images. The images in the gallery and the probe sets are of low resolution. This setting and the natural LQ property make it a challenging dataset. For the third category, we choose IJB-B [39] as the mixed-quality dataset which contains about 1.8K identities with a total of 21.8K images and 55K unconstrained video frames. In the 1:1 verification task, there are about 10K positive matches and 8M negative matches. Examples of these datasets are shown in Fig. 4.
@@ IV-B Ablation on different training strategies
Setting A: To show the influence of the LQ data, we apply the data augmentations recommended in AdaFace, i.e. down-sampling with a following up-sampling, cropping with a small part, and random color distortion. All these three augmentations are applied with independent possibilities of 0.2.
Setting B: To show the negative influence of some extreme LQ data on the classification-based method, we apply quality partitioning on AdaFace loss by simply ignoring them (the contrastive learning is not presented).
@@
Setting D: To show the impacts of data augmentation, we apply our augmentation with the original AdaFace.
@@
Table I shows the results of the above settings. The Baseline setting gains solid performances on the HQ datasets and IJB-B dataset. However, its performances on d1 of SCface dataset and Tinyface dataset are significantly poor. In setting A, when data augmentation empowers the original feature extractor to gain a higher performance on the HQ and the LQ datasets, the accuracy dramatically drops on the IJB-B dataset. This could be interpreted as a sacrifice of the discriminability. With a single classification supervision signal, the features become less distinctive to compromise between the HQ and the LQ images. On IJB-B dataset where the mixed-quality matching between images requires stronger discriminability, this feature degradation misleads the comparisons among a large number of faces with various qualities. Setting B shows that only dropping some LQ data helps the model gain better understanding on the LQ SCface dataset. This illustrates that some extremely low-quality data could mislead the training of the classification method. Even though AdaFace adapts the data importance according to the quality, it is harmful to learn these LQ data with classification loss. In Setting C, our approach relieves the degradation caused by data augmentation on IJB-B and outperforms the Baseline settings. Furthermore, the performance on SCface dataset is also improved. Meanwhile, our performance on the HQ data is only slightly dropped by less than 0.05% compared with Setting B. This shows the effectiveness of contrastive learning on the LQ data and the ability of our method to keep the comprehension on the HQ data. Setting D improves the performances on the LQ datasets and demonstrates the effectiveness of our data augmentation. Meanwhile, it is inferior to Setting C on IJB-B dataset which illustrates the effectiveness of our joint training framework again.
@@
Finally, we implement Setting E, which is our proposed approach. AdaFace learns the hard-to-learn LQ images by classification loss with relatively less importance. Instead, we learn from them with contrastive learning. In this way, our model keeps learning the LQ images and gains more instance-level understanding from these less discriminative images. As a result, our method achieves the best results on almost all the evaluation datasets of different qualities. To ensure that most of the augmented LQ images are discriminable and realistic, we down-sample each image and then up-sample it to the original resolution. The discarding ratio, the degree of the brightness variations, and the contrast and saturation in color distortion are set to lower levels. In addition, we follow C. Kuo et al. [6] to add JPEG compression artifacts to make the image more realistic. With the new data augmentation, the augmented images become more understandable and compliant with the real-world low-resolution images. Fig. 5 shows some examples. The image pairs in red box are marked as LQ. They are harder to be identified than others in the training data (in green box) due to the blur, the obstacles, or the non-frontal positions in the images. Our quality partitioning strategy sends these LQ data from the original and the augmented images to contrastive learning, and this setting outperforms all the other settings on SCface and Tinyface. Furthermore, it also shows promising results on the IJB-B dataset which contains a huge number of image pairs of mixed qualities.
@@ IV-C Ablation on the contrastive queue
TABLE II: Ablation of different queue settings. “No” indicates applying data augmentation and quality partitioning without contrastive loss. “Batch” indicates applying our quality-guided training strategy with the samples in batch as the negative pool for contrastive learning.
Queue SCface Tinyface IJB-B HQ Avg. LFW CFP-FP AgeDB CPLFW CALFW
@@
No 87.85 99.38 100.0 68.24 90.57 95.47 99.45 96.00 94.20 94.67 93.05
Batch 87.54 99.38 100.0 68.67 87.38 95.28 99.55 95.96 94.38 93.85 92.68
Momentum 63.08 97.85 99.85 62.34 88.70 95.18 99.55 96.03 93.98 94.65 91.68
Proxy-updated 92.31 99.54 100.0 69.85 91.05 95.43 99.50 96.03 94.00 94.48 93.12
Fig. 3 shows the average difference of the feature pairs with the momentum queue and our proxy-updated real-time queue. As the momentum queue is constructed with the features from a slowly-progressing momentum encoder, the features are steady and whitened. Both the positive and the negative pairs are limited in a tight boundary. While in real scenarios where the difference is computed with the features from the same training encoder, it shows a much larger gap between the query and the corresponding key. Our proxy-updated queue takes such features as the positive pairs in contrastive learning directly and updates the features in queue with the identity proxies. According to the large size of the queue compared with the training batch, the difference between the negative pairs shows smaller magnitudes compared with the positive features in our queue.
Table II shows the performances of the models with different queue settings. In the “No” situation, contrastive learning is not applied, while both the original and the augmented images are used for learning. In the “Batch” setting, the queue size is limited to the current batch. The number of feature-feature pairs in contrastive learning is extremely smaller than the feature-identity pairs in the classification loss. Thus, it stumbles on the IJB-B dataset which contains a large number of mixed-quality image pairs and needs higher discriminability. The momentum queue utilizes the whitened features to compose the positive and the negative pairs in contrastive learning. This misleads the optimization process in our joint training scenario. As a result, the performances of all datasets are lower than the setting without a queue. Our proxy-updated queue provides genuine positive pairs and a dynamic negative queue to keep the same magnitude on the number of pair similarity computation as in the classification process. This enables the contrastive loss to effectively learn from the LQ images and keep high discriminability. As a result, it outperforms the baseline and the other settings equipped with other queues on all benchmarks except a slight performance drop ($0.04\%$) on the HQ datasets.
@@ IV-D Ablation on the quality partitioning threshold
Threshold SCface Tinyface IJB-B HQ Avg. LFW CFP-FP AgeDB CPLFW CALFW
@@ IV-E Analysis on the SCface dataset
Besides the quantitative analysis above, we further illustrate the difference between our QGFace and the Baseline on SCface in Fig.6. It is obvious that the matched pairs encoded by our QGFace are more similar than the Baseline method, which demonstrates the effectiveness of the quality partitioning strategy, the contrastive learning, and the selected data augmentation. As the evaluation process aims at finding an image among all of the candidates in gallery, the most similar unmatched image is also important for the successful match. We then calculate the statistics of the similarity between a probe and its matched gallery image, and the similarity between the probe and its most similar unmatched gallery image. The difference between the means of the two kinds of similarity scores is calculated, and the improvement of our method is noted in the title of the sub-figures. When the distances become longer and the qualities of the images become lower, the improvement of QGFace gets larger. In the case of 4.2m, the improvement is even more than 600%, which clearly explains how our method improves the accuracy by more than 30% on SCface-d1 in Table I.
@@ IV-F Comparison with SOTA methods
On SCface dataset. Table IV compares our method with the SOTA methods on SCface dataset. Many quality-invariant methods need a fine-tuning process to learn the target domain information. This breaks the general end-to-end framework and limits the evaluation confidence as only 130 identities and no distracting faces are provided in SCface. Furthermore, all these methods show lower accuracy scores in d3 compared with the pure AdaFace trained with the HQ data. AdaFace-Aug, which is the same as Setting A in Table I, achieves 100.0% by learning the relatively-high-quality augmented images. This verifies our hypothesis that the quality-invariant methods lose discriminative information during training. Our approach achieves the best performances on all the three distances with clear margins compared with the other approaches. This demonstrates the effectiveness of our proposed quality-guided joint training strategy for mixed-quality face recognition by simultaneously learning both the HQ and the LQ images.
On Tinyface dataset. Table V compares our method with other SOTA methods on Tinyface dataset. For the pure classification-based CurricularFace, even incorporated with a larger backbone model and a larger training dataset, it does not show an outstanding result compared with other methods. AdaFace puts relatively less emphasis on the hard LQ images, and outperforms CurricularFace. However, it is still under-competitive with the methods that focus on LQ face recognition. Our method partitions the data according to the image qualities and learns the LQ images with contrastive learning instead of harder classification process. The approach in RIFR [11] slightly outperforms our method by employing a distillation approach. This needs an extra pre-trained HQ face encoder to transfer its knowledge to the trained encoder, while our method does not have this preliminary requirement and is orthogonal with it. As we focus on training a feature extractor that is capable of dealing with the mixed-quality images, it is easy to integrate QGFace with other methods to further improve the performance. Furthermore, our method achieves a better average performance on SCface and Tinyface compared with RIFR [11], which demonstrates the effectiveness of our QGFace in recognizing mixed-quality face images under different scenarios.
TABLE IV: Rank-1 accuracies (%) of different methods on SCface. “FT” indicates that the method is fine-tuned with a part of the SCface dataset. “AdaFace-Aug” is the same as setting A in Table I.
@@
Synthesis SR-DCR (ICCPR20) [27] 74.10 93.70 97.2
@@
Encoder- Only NPT-Loss (TPAMI22) [46] 85.69 99.08 99.08

### v2

@@ III-B2 Contrastive learning for LQ features
Although classification-based methods achieve SOTA performances in HQ face recognition, they are not suitable for learning LQ images. Fig. 2 visualizes the GST of AdaFace. It is obvious that LQ images gain more attention during training. However, due to the low qualities, they lack the discriminative information for matching with thousands of abstractive identity features. To alleviate this problem, we apply contrastive learning on LQ features, which is based on the similarity between image instances and is more practical for learning the LQ data.
@@
To comprehensively learn the LQ images and compose the positive contrastive pairs, we apply data augmentation on the HQ training images. Unlike popular contrastive methods that apply data augmentation on images twice to compose the positive pairs, we only augment each image once to compose the pair with the original image. In addition, we stop the gradients of the HQ images during training to keep the qualities of their features. Let $(\boldsymbol{x}^{q},\boldsymbol{x}^{k_{+}})$ denote the augmented image and the original image, respectively, and $(\boldsymbol{h}^{q},\boldsymbol{h}^{k_{+}})$ denote their features, which compose a positive pair in contrastive learning.
Different augmentation methods are used, including scaling (downsampling to a lower resolution and then upsampling to the original size), crop & random resize, rotation, and color distortion. AdaFace introduces hard data augmentations with low possibilities including cropping 20% of the original image, which makes the augmented images unrecognizable. This might be acceptable when little emphasis is put on these data, but is not suitable in our situation where all LQ features are used for learning. We increase the possibility of data augmentation with lower-level distortion. In addition, we follow [19] to add JPEG compression artifacts to make the augmented images more realistic.
Finally, to better cooperate with the above two learning paradigms, we employ the Supervised Contrastive Mask (SCM) introduced in CoReFace [36]. Pair-guided contrastive learning may pull two features away as they are not generated from the same image, while the classification method tries to pull them together as they are from the same identity. To address this conflict, SCM removes all samples with the same identity of the given image from the negative feature pool. Let $\theta_{qk}$, $y_{+}$ denote the angle between feature pairs and the label of the given image, respectively, and $Q$, $k_{j}$ denote the number of negative samples and the $j^{th}$ feature among them. With a scale parameter $s$, the contrastive loss can be formulated as
@@ III-C Proxy-Updated Real-Time Queue
Self-supervised contrastive learning usually needs a large comparison pool to serve as negative candidates and push the training samples away from them. MoCo [8] defines the self-supervised contrastive learning as the comparison with a queue. A slowly progressing momentum encoder is proposed to produce keys and save them in a queue. The queries are produced by the training encoder. To make the positive and negative keys comparable, all of them are produced by the momentum encoder. However, this operation prevents the training encoder from learning the LQ and HQ images simultaneously, as one of them is sent into a gradient-free encoder. Furthermore, as the momentum encoder is designed to be stable during large training intervals, this could lose track of the status of the encoder in joint training and produce whitened features.
To relieve the above problems, we propose a novel proxy-updated queue for contrastive learning, which is illustrated in Fig. 3. We take the genuine features encoded by the training model as positive pairs. Since the training target of the cross-entropy softmax loss could constrain the corresponding feature-proxy pair to be similar, we update the negative features with their related proxies. Thus, we build a feature queue and a proxy queue. Specifically, we first encode all the original images and the augmented ones. Then, we update the feature queue and the proxy queue with the above features and their corresponding identity proxies. As the features in queue could be generated from several iterations ago, we update them to be similar to the features generated by the training encoder. They are added with the difference between the proxies in queue and the proxies in training. To provide a mixed-quality queue, we take both $\boldsymbol{h}^{q}$ and $\boldsymbol{h}^{k}$ to update it. With SCM, no comparisons between identical features will be introduced.
The cross-entropy softmax loss calculates the similarity between a feature and all identity proxies. To provide a comparative magnitude of pair relationship supervision, we set the size of the queue $Q$ as the number of the identities $n$. Thus, the number of pair comparisons used for contrastive learning keeps growing in training. This enables contrastive learning to discriminate among more images of mixed qualities for more effective learning.
@@ IV Experiments
[…]
Setting
Augmentation
Quality Contrastive SCface
Tinyface
IJB-B
HQ Avg.
@@ IV-A Datasets and Implementation Details
Datasets. We take VGGFace2 [2] as the training dataset. We evaluate our approach on three categories of datasets for different tasks. The first category is the HQ datasets including widely used LFW [10], AgeDB [30], CFP-FP [33], CPLFW [50], and CALFW [51]. The second category is the LQ datasets including the SCface dataset [7] and Tinyface dataset [4] which are usually used for cross-resolution face recognition and LQ face recognition, respectively. SCface contains 130 identities captured by five surveillance cameras at 3 distances: 4.2 m (d1), 2.6 m (d2) and 1.0 m (d3) for each identity. At each distance, the LQ images need to be matched with the HQ gallery images of the corresponding identities. Tinyface contains 5K identities with 153.5K distracting images. The images in the gallery and the probe sets are of low resolution. This setting and the natural LQ property make it a challenging dataset. For the third category, we choose IJB-B [43] as the mixed-quality dataset which contains about 1.8K identities with a total of 21.8K images and 55K unconstrained video frames. In the 1:1 verification task, there are about 10K positive matches and 8M negative matches.
@@ IV-B Ablation on different training strategies
Setting A: To show the influence of the LQ data, we apply the data augmentations recommended in AdaFace, i.e. down-sampling with a subsequent up-sampling, cropping with a small part, and random color distortion. All these three augmentations are applied with independent possibilities of 0.2.
Setting B: To show the negative influence of some extreme LQ data on the classification-based method, we apply quality partitioning on AdaFace loss simply by ignoring them (contrast learning is not presented).
@@
Setting D: To show the impacts of data augmentation, we apply our augmentation to the original AdaFace.
@@
Table I shows the results of the above settings. The baseline setting gains solid performance on the HQ datasets and IJB-B dataset. However, its performances on d1 of SCface dataset and Tinyface dataset are significantly poor. In setting A, when data augmentation empowers the original feature extractor to gain higher performance on the HQ and the LQ datasets, the accuracy dramatically drops on the IJB-B dataset. This could be interpreted as a sacrifice of discriminability. With a single classification supervision signal, the features become less distinctive in a compromise between the HQ and the LQ images. On the IJB-B dataset where mixed quality matching between images requires stronger discriminability, this feature degradation misleads the comparisons among a large number of faces with various qualities. Setting B shows that only dropping some LQ data helps the model gain better understanding on the LQ SCface dataset. This illustrates that some extremely low-quality data could mislead the classification method training. Even though AdaFace adapts the data importance according to the quality, it is harmful to learn these LQ data with classification loss. In Setting C, our approach relieves the degradation caused by data augmentation in IJB-B and outperforms the baseline settings. Furthermore, the performance on SCface dataset is also improved. Meanwhile, our performance on the HQ data is only slightly dropped by less than 0.05% compared with Setting B. This shows the effectiveness of contrastive learning on the LQ data and the ability of our method to keep the comprehension on the HQ data. Setting D improves the performance on the LQ datasets and demonstrates the effectiveness of our data augmentation. Meanwhile, it is inferior to Setting C on IJB-B dataset which illustrates the effectiveness of our joint training framework again.
@@
Finally, we implement Setting E, which is our proposed approach. AdaFace learns hard-to-learn LQ images by classification loss with relatively less importance. Instead, we learn from them with contrastive learning. In this way, our model continues to learn the LQ images and gains more instance-level understanding from these less discriminative images. As a result, our method achieves the best results on almost all evaluation datasets of different qualities. To ensure that most of the augmented LQ images are discriminable and realistic, we down-sample each image and then up-sample it to the original resolution. The discard ratio, the degree of brightness variations, and the contrast and saturation in color distortion are set to lower levels. In addition, we follow C. Kuo et al. [19] to add JPEG compression artifacts to make the image more realistic. With the new data augmentation, the augmented images become more understandable and compliant with the real-world low-resolution images. Fig. 4 shows some examples. The image pairs in the red box are marked as LQ. They are harder to be identified than others in the training data (in green box) due to the blur, the obstacles, or the non-frontal positions in the images. Our quality partitioning strategy sends these LQ data from the original and the augmented images to contrastive learning, and this setting outperforms all the other settings on SCface and Tinyface. Furthermore, it also shows promising results on the IJB-B dataset which contains a huge number of image pairs of mixed qualities.
@@ IV-C Ablation on the contrastive queue
TABLE II: Ablation of different queue settings. “No” indicates no contrastive loss. “Batch” indicates the samples in batch are taken as the negative pool.
Queue
SCface
Tinyface
IJB-B
HQ Avg.
@@
No 87.85 99.38 100.0 68.24 90.57 95.47
Batch 87.54 99.38 100.0 68.67 87.38 95.28
Momentum 63.08 97.85 99.85 62.34 88.70 95.18
Proxy-updated 92.31 99.54 100.0 69.85 91.05 95.43
Fig. 3 shows the average difference of the feature pairs with the momentum queue and our proxy-updated real-time queue. As the momentum queue is constructed with the features of a slowly progressing momentum encoder, the features are steady and whitened. Both the positive and the negative pairs are limited in a tight boundary. While in real scenarios where the difference is computed with the features from the same training encoder, it shows a much larger gap between the query and the corresponding key. Our proxy-updated queue takes these features as the positive pairs in contrastive learning directly and updates the features in queue with the identity proxies. According to the large size of the queue compared with the training batch, the difference between the negative pairs shows smaller magnitudes compared with the positive features in our queue.
Table II shows the performance of the models with different queue settings. In the “No” situation, contrastive learning is not applied, while the original and the augmented images are used for learning. In the “Batch” setting, the queue size is limited to the current batch. The number of feature-feature pairs in contrastive learning is extremely smaller than the feature-identity pairs in the classification loss. Thus, it stumbles on the IJB-B dataset which contains a large number of mixed-quality image pairs and needs higher discriminability. The momentum queue utilizes whitened features to compose positive and negative pairs in contrastive learning. This misleads the optimization process in our joint training scenario. As a result, the performances of all datasets are lower than the setting without a queue. Our proxy-updated queue provides genuine positive pairs and a dynamic negative queue to keep the same magnitude on the number of pair similarity computation as in the classification process. This enables contrastive loss to effectively learn from LQ images and maintain high discriminability. As a result, it outperforms the baseline and the other settings equipped with other queues on all benchmarks except a slight performance drop ($0.04\%$) on the HQ datasets.
@@ IV-D Ablation on the quality partitioning threshold
Threshold
SCface
Tinyface
IJB-B
HQ Avg.
@@ IV-E Analysis on the SCface dataset
Besides the quantitative analysis above, we further illustrate the difference between our QGFace and the Baseline on SCface in Fig.5. It is obvious that the matched pairs encoded by our QGFace are more similar than the Baseline method, which demonstrates the effectiveness of the quality partitioning strategy, the contrastive learning, and the selected data augmentation. As the evaluation process aims to find an image among all of the candidates in the gallery, the most similar unmatched image is also important for the successful match. We then calculate the statistics of the similarity between a probe and its matched gallery image, and the similarity between the probe and its most similar unmatched gallery image. The difference between the means of the two kinds of similarity scores is calculated, and the improvement of our method is noted in the title of the sub-figures. When the distances become longer and the qualities of the images become lower, the improvement of QGFace becomes larger. In the case of 4.2m, the improvement is even more than 600%, which clearly explains how our method improves the accuracy by more than 30% on SCface-d1 in Table I.
@@ IV-F Comparison with SOTA methods
On SCface dataset. Table IV compares our method with SOTA methods in SCface dataset. Many quality-invariant methods need a fine-tuning process to learn the target domain information. This breaks the general end-to-end framework and limits the evaluation confidence as only 130 identities and no distracting faces are provided in SCface. Furthermore, all these methods show lower accuracy scores in d3 compared with the pure AdaFace trained with the HQ data. AdaFace-Aug, which is the same as Setting A in Table I, achieves 100.0% by learning the relatively high quality augmented images. This verifies our hypothesis that the quality-invariant methods lose discriminative information during training. Our approach achieves the best performance on all three distances with clear margins compared with the other approaches. This demonstrates the effectiveness of our proposed quality-guided joint training strategy for mixed-quality face recognition by simultaneously learning both the HQ and the LQ images.
On Tinyface dataset. Table V compares our method with other SOTA methods on Tinyface dataset. For the pure classification-based CurricularFace, even incorporated with a larger backbone model and a larger training dataset, it does not show an outstanding result compared with other methods. AdaFace puts relatively less emphasis on the hard LQ images and outperforms CurricularFace. However, it is still under-competitive with the methods that focus on LQ face recognition. Our method partitions the data according to the image qualities and learns the LQ images with contrastive learning instead of harder classification process. The approach in RIFR [14] slightly outperforms our method by employing a distillation approach. This needs an extra pre-trained HQ face encoder to transfer its knowledge to the trained encoder, while our method does not have this preliminary requirement and is orthogonal with it. As we focus on training a feature extractor capable of dealing with mixed-quality images, it is easy to integrate QGFace with other methods to further improve performance. Furthermore, our method achieves a better average performance on SCface and Tinyface compared with RIFR [14], which demonstrates the effectiveness of our QGFace in recognizing mixed-quality face images under different scenarios.
TABLE IV: Rank-1 accuracies (%) of different methods on SCface. “FT” indicates that the method is fine-tuned. “AdaFace-Aug” is the same as setting A in Table I.
@@
W. Hainan et al. (FG23) [40] 89.00 99.00 99.25
CATFace (TBIOM24) [37] 90.64 98.85 99.61
Synthesis
SR-DCR (ICCPR20) [32] 74.10 93.70 97.2
@@
Encoder-
Only
NPT-Loss (TPAMI22) [15] 85.69 99.08 99.08